Solve system using the subsitution method"
x^2+3y=25
x+y=7

Answer by 24HoursTutor.com(40)

x^2+3y = 25
and
x+y=7
So, y = 7 - x
Now, we substitute this value of y in the first equation to get :
x^2 + 3(7-x) = 25
x^2 + 21 - 3x = 25
x^2 -3x= 25 - 21
x(x-3) = 4 ( We take the common multiple x outside )
x = 4
Since x = 4
y = 7- x = 7-4 = 3
Ans. x=4, y=3
We can also check our answers by substituting the correct values in the given equations :
x^2+3y=25
4^2+3X3 = 25
16 + 9 = 25
Which means our values for x and y are correct.
The above question is solved by one of the experts from 24HoursTutor.com
We provide live 1:1 online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr